OpenStudy (anonymous):

What is the slope of the line that passes through the points (2, 3) and (–2, 6)? A.3/4 B.0 C.-3/2 D.-3/4

OpenStudy (anonymous):

@selfconfidence777 need help in math

OpenStudy (selfconfidence777):

b.

OpenStudy (anonymous):

Which is the equation of the line? A. 4x + y = 2 B. x + 4y = –8 C. 4x – y = –2 D. x – 4y = –8

OpenStudy (anonymous):

@selfconfidence777

OpenStudy (selfconfidence777):

A

OpenStudy (selfconfidence777):

or C

OpenStudy (selfconfidence777):

not sire but C looks good ,

OpenStudy (anonymous):

GREAT MY COMPUTER JUST REBOOTED MY TEST